OpenStudy (anonymous):

-(7/5u) -(3/5) = 7u -(1/2) Simplified as much as possible. I think the answer is x = 85

OpenStudy (anonymous):

is the \(u\) in the denominator?

OpenStudy (anonymous):

It's really weird to write...let me use the equations this program offers and rewrite it

OpenStudy (anonymous):

\[-\frac{ 7 }{ 5 } u -\frac{ 3 }{ 5 } = 7u - \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

mutiply by 10 to clear the annoying fractions and start with \[-14x-6=70x-5\]

OpenStudy (anonymous):

-20x = 65x?

OpenStudy (anonymous):

and then simplify that?

OpenStudy (anonymous):

you lost me

OpenStudy (anonymous):

\[-14x-6=70x-5\] a) add \(14x\) to both sides b) add 5 to both sides c) divide both sides by \(86\)

OpenStudy (anonymous):

oh duh! I forgot to get the x by itself...rookie mistake

OpenStudy (anonymous):

how did you get the 86? wouldn't it be 89?

OpenStudy (anonymous):

lol it would be neither \(70+14=84\)!!

OpenStudy (anonymous):

so x-1 = 84x?

OpenStudy (anonymous):

I'm a bit stuck